# Advanced Vision Transformers and Open-Set Learning for Robust Mosquito Classification: A Novel Approach to Entomological Studies

### v1

@@ 4 Results and Discussion
This section represents the findings of this study for both closed and open-set learning for static and non-stationary environments, respectively. The first subsection contains mosquito image classification for closed-set. Next, the open-set learning results are illustrated for the applied deep learning techniques employing the OpenMax dataset and Weibull distribution. Finally, the proposed mosquito classification system has been compared with the related articles.
@@ 4.1 Results for Closed-set Learning
Table 2 presents the performance metrics for various models applied to closed-set learning in the proposed mosquito classification. The Swin-S Transformer technique achieves the highest overall accuracy at 99.80%, with corresponding high precision, recall, and F1 scores, indicating its superior performance. MobileViT achieves satisfactory performance with a high accuracy of 98.90% and an F1 score of 0.989, while utilizing significantly fewer parameters (5.58 million), making it an efficient model for mosquito classification. Conversely, ResNet50 shows the lowest accuracy at 94.50%, but still demonstrates reliable performance with a precision of 0.937 and a recall of 0.933. These results highlight the effectiveness of transformer-based models, particularly Swin-S and Swin-B, in accurately classifying mosquito species.
Model Precision Recall F1 score Accuracy () Parameters (million) ResNet50 0.937 0.933 0.933 94.50 23.59 ViT Base 0.979 0.978 0.977 97.80 85.81 CvT-13 0.985 0.984 0.984 98.40 19.62 MobileViT 0.989 0.989 0.989 98.90 5.58 Xception 0.991 0.911 0.991 99.20 22.91 Swin-B 1.000 1.000 1.000 99.70 86.75 Swin-S 0.998 0.998 0.998 99.80 48.84
@@
Figure 8 represents the normalized confusion matrix of the applied models for the test samples of the ten mosquito classes. According to Figure 8(a), the Swin Transformer model demonstrates high accuracy, with most species correctly classified. However, there are minor misclassifications, such as Anopheles arabiensis being confused with Anopheles Coluzzii and Anopheles Atroparvus. The MobileViT model attained almost similar performance with Swin in mosquito classification. This lightweight model with significantly reduced parameters achieves high classification accuracy overall, although some misclassifications are observed, such as Anopheles Arabiensis being occasionally misidentified as Anopheles coluzzii. The ViT Transformer model shows exceptional performance, though some misclassifications, such as Anopheles arabiensis being occasionally misidentified as Anopheles coluzzii and Anopheles albimanus.
Figure 9 illustrates the ROC curves for different models applied to closed-set learning, including Swin-B, Xception, ViT, and Swin-S. The curves demonstrate the high performance of the applied models, with most achieving near-perfect AUC scores, indicating their strong capability in distinguishing between mosquito species.
@@ 4.2 Open-set Learning Implementation Results
The primary challenge we encountered during our research was that our models were susceptible to misclassification when presented with images outside the domain of our dataset. Our models, which were trained exclusively on our mosquito dataset, incorrectly classified images of unknown classes (e.g., ants, bees, butterflies) as one of the ten known mosquito classes. This issue is illustrated in Figure 11, where even images of ants, bees, and butterflies were erroneously predicted to belong to one of the mosquito classes. Figure 11(a) shows that an ant was classified as Aedes albopictus with a confidence score of 0.54. In figure 11(b),11(c), two different types of bee were classified as Aedes aegypti and Culex quinquefasciatus with scores of 0.47 and 0.87, respectively. In figure 11(d),11(e), two different types of butterfly were classified as Aedes aegypti with scores of 0.77 and 0.71, respectively. To address this problem, we incorporated the OpenMax method, which allowed us to effectively identify and label outlier images as unknown, thereby improving the robustness of our model in handling images from outside the training dataset.
The primary challenge encountered in our research was the susceptibility of the applied Swin model to misclassifications when presented with images outside the domain of our dataset. Our models, trained exclusively on mosquito images, incorrectly classified unknown insects, such as ants, bees, and butterflies, as one of the known mosquito classes, as illustrated in Figure 11. These misclassifications are expected in a static environment, as the applied models attained incomplete knowledge during training, leading to reduced robustness and flexibility. For instance, an ant was classified as Aedes albopictus with a confidence score of 0.54, while bees and butterflies were erroneously predicted as various mosquito species. To address this issue, we implemented the open-set learning technique with the OpenMax method, which effectively identified and labeled outlier images as unknown. This enhancement improved the robustness of our model, allowing it to better handle images from outside the training dataset and reducing the misclassification rate.
@@
Study Dataset Number of Classes Best Model Accuracy (%) F1 Score [6] IEEE data port 2 ETC 99.20% 0.99 [7] Open-source, 2400 images 2 VGG16 98.25% 0.95 [8] Custom, 5400 images 6 VGG16 91.17% N/A [11] 9900 images 7 Swin Transformer 96.30% 0.98 [12] Custom, 397 images 2 QSVM 93% N/A [15] N/A 16 Xception 97.040.87% 0.97 [16] 2648 images 3 GoogLeNet 96% N/A Proposed Merged 4 public datasets 5000 images 10 Swin-B Transformer 99.70% 1.0
Table 4 compares the proposed mosquito classification system with related studies. The proposed system, which merges four public datasets totaling 5000 images and ten mosquito classes and uses the Swin-B model, achieves the highest accuracy of 99.7%.

### v2

@@ 4 Results and Discussion
This section represents the findings of this study for both closed and open-set learning for static and non-stationary environments, respectively. The first subsection contains mosquito image classification for closed-set. All the evaluations of the model’s performance were performed using the unseen data during training or test set. Next, the open-set learning results are illustrated for the applied deep learning techniques employing the OpenMax dataset and Weibull distribution using the extended test set data, specially prepared for this particular task. Finally, the proposed mosquito classification system has been compared with the related articles.
@@ 4.1 Results for Closed-set Learning
Table 3 presents the performance metrics for various models applied to closed-set learning in the proposed mosquito classification. The Swin-B Transformer achieves the highest overall accuracy at 99.60%, with corresponding high precision (0.994), recall (0.996), and F1 score (0.996), indicating its superior performance. The MobileViT model also performs well, achieving a high accuracy of 98.90% and an F1 score of 0.989, while utilizing significantly fewer parameters (5.58 million), making it an efficient model for mosquito classification. The Xception model follows closely with an accuracy of 99.20% and an F1 score of 0.991. On the other hand, ResNet50 shows the lowest accuracy at 94.50%, but still demonstrates reliable performance with a precision of 0.937 and a recall of 0.933. All the results were based on the fine-tuned model’s performance on unseen data during the training phase, the test set. These results highlight the effectiveness of transformer-based models, particularly Swin-B and MobileViT, in accurately classifying mosquito species while maintaining computational efficiency.
Model Precision Recall F1 score Accuracy () Parameters (million) ResNet50 0.937 0.933 0.933 94.50 23.59 ViT Base 0.979 0.978 0.977 97.80 85.81 CvT-13 0.985 0.984 0.984 98.40 19.62 MobileViT 0.989 0.989 0.989 98.90 5.58 Xception 0.991 0.911 0.991 99.20 22.91 Swin-B 0.996 0.996 0.996 99.60 86.75 Swin-S 0.983 0.981 0.981 98.25 48.84
@@
Figure 8(d) represents the normalized confusion matrix of the applied models for the test set samples of the ten mosquito classes (unseen data during the training phase). According to Figure 8(a), the Swin Transformer model demonstrates high accuracy, with most species correctly classified. However, there are minor misclassifications, such as Anopheles arabiensis being confused with Anopheles Coluzzii and Anopheles Atroparvus. The MobileViT model attained almost similar performance with Swin in mosquito classification. This lightweight model with significantly reduced parameters achieves high classification accuracy overall, although some misclassifications are observed, such as Anopheles Arabiensis being occasionally misidentified as Anopheles coluzzii. The ViT Transformer model shows exceptional performance, though some misclassifications, such as Anopheles arabiensis being occasionally misidentified as Anopheles coluzzii and Anopheles albimanus.
Figure 9(d) illustrates the ROC curves for different models applied to closed-set learning, including Swin-B, Xception, ViT, and Swin-S. The curves demonstrate the high performance of the applied models, with most achieving near-perfect AUC scores, indicating their strong capability in distinguishing between mosquito species.
@@
The Swin-B model, known for its high precision in classifying individual mosquitoes, also excels at detecting and categorizing multiple mosquitoes within a single image. Thanks to its transformer-based architecture, which captures intricate spatial relationships and minute details, the model can effectively isolate and recognize each mosquito, even in complex or cluttered environments. By employing attention mechanisms, Swin-B can focus on various sections of the image, allowing it to detect and classify multiple mosquitoes, despite challenges like overlapping or different poses. This ability makes Swin-B particularly valuable for real-world applications where mosquitoes often appear in groups.
As shown in Figure 11(a), Figure 11(b) and Figure 11(c), the model demonstrates its capability to detect mosquitoes accurately in images where multiple mosquitoes are present, identifying each mosquito individually even in crowded scenes.
@@ 4.2 Open-set Learning Implementation Results
The primary challenge we encountered during our research was that our models were susceptible to misclassification when presented with images outside the domain of our dataset. Our models, which were trained exclusively on our mosquito dataset, incorrectly classified images of unknown classes (e.g., ants, bees, butterflies) as one of the ten known mosquito classes. This issue is illustrated in Figure 12, where even images of ants, bees, and butterflies were erroneously predicted to belong to one of the mosquito classes. Figure 12(a) shows that an ant was classified as Aedes albopictus with a confidence score of 0.54. In figure 12(b),12(c), two different types of bee were classified as Aedes aegypti and Culex quinquefasciatus with scores of 0.47 and 0.87, respectively. In figure 12(d),12(e), two different types of butterfly were classified as Aedes aegypti with scores of 0.77 and 0.71, respectively. To address this problem, we incorporated the OpenMax method, which allowed us to effectively identify and label outlier images as unknown, thereby improving the robustness of our model in handling images from outside the training dataset.
@@
Table 5 compares the proposed mosquito classification system with related studies. The proposed system, which merges four public datasets totaling 5000 images and ten mosquito classes and uses the Swin-B model, achieves the highest accuracy of 99.6%.
Study Dataset and Number of Images After Augmentation Number of Classes Best Model Accuracy (%) F1 Score [6] IEEE data port 2 ETC 99.20% 0.99 [7] Open-source, 2400 images 2 VGG16 98.25% 0.95 [8] Custom, 5400 images 6 VGG16 91.17% N/A [11] 9900 images 7 Swin Transformer 96.30% 0.98 [12] Custom, 397 images 2 QSVM 93% N/A [15] N/A 16 Xception 97.040.87% 0.97 [16] 2648 images 3 GoogLeNet 96% N/A Proposed Merged 4 public datasets 5000 images 10 Swin-B Transformer 99.60% 0.996